# Quark matter, weak interactions, magnetic field, Landau levels, neutrino emission, quark stars, pulsar kicks

### v1

@@ 3.1 Neutrino number production rate
By substituting the Green functions given in Eqs. (5) and (6) into Eq. (46) and neglecting neutrino trapping (i.e., assuming that the neutrino and antineutrino distribution functions vanish), we obtain the following expression for the self-energies:
@@ 4 Summary and Outlook
As an outlook for future work, two directions warrant further investigation. First, color-superconducting phases of quark matter should be explored, as diquark pairing can modify quasiparticle spectra, level occupancies, and gap structures, potentially enhancing anisotropic emission and momentum asymmetries. Second, the early deleptonization stage of compact stars, in which neutrino trapping and diffusion are essential, has not been addressed in the present study. In this regime, anisotropic neutrino diffusion in a strongly magnetized, lepton-rich medium may generate momentum asymmetries qualitatively different from those found in the free-streaming limit considered here. Both directions require detailed microscopic calculations and fully self-consistent transport modeling.

### v2

@@ 3.1 Neutrino number production rate
By substituting the Green’s functions given in Eqs. (8) and (9) into Eq. (49) and neglecting neutrino trapping (i.e., assuming that the neutrino and antineutrino distribution functions vanish), we obtain the following expression for the self-energies:
@@ 4 Summary and Outlook
As an outlook for future work, at least two directions warrant further investigation. First, the early deleptonization stage of compact stars, in which neutrino trapping and diffusion are essential, has not been addressed in the present study. In this regime, anisotropic neutrino diffusion in a strongly magnetized, lepton-rich medium may generate momentum asymmetries qualitatively different from those found in the free-streaming limit considered here. This requires detailed microscopic calculations at nonzero neutrino chemical potential and self-consistent transport modeling.
Second, since quark matter is expected to form Cooper pairs and enter a color-superconducting state at sufficiently high densities Shovkovy (2005); Alford et al. (2008), it is important to investigate the impact of diquark pairing on neutrino emission and the associated momentum asymmetries. Without detailed calculations, one can argue that the results should depend sensitively on the underlying diquark pairing pattern Schäfer and Schwenzer (2004).
In the two-flavor color-superconducting (2SC) phase, a third of quark modes remain ungapped, so its emission should resemble that of normal quark matter, with the rate about three times smaller Jaikumar et al. (2006); Alford et al. (2025). Indeed the contribution of gapped modes is exponentially suppressed at low temperatures, scaling as $\dot{\cal E}_{\nu}\propto e^{-\Delta/T}$, where $\Delta\sim\mathcal{O}(10\,\mathrm{MeV})$ denotes the color-superconducting gap. At sufficiently low temperatures, such modes therefore contribute very little.
By contrast, in the color-flavor-locked (CFL) phase all quark modes are gapped, strongly suppressing low-energy excitations and rendering negligible neutrino emission at low temperatures. Notably, the specific heat of CFL phase is also exponentially suppressed and, thus, the cooling of compact stars with quark cores in the CFL phase is a delicate issue Shovkovy and Ellis (2002).
Finally, single-flavor spin-one color-superconducting phases admit multiple possible pairing patterns, each requiring separate analysis. Of particular interest are those phases in which the gap exhibits nodes or lines on the Fermi surface. Such gapless phases permit low-energy excitations and can therefore lead to non-negligible neutrino emission with potentially nontrivial consequences for stellar cooling Schmitt et al. (2006).
These general arguments established for color superconductors in the zero-field limit may remain qualitatively valid in the presence of weak magnetic fields. However, the interplay between color superconductivity and strong magnetic fields is far from fully understood. Intense magnetic fields can alter the gap structure Ferrer et al. (2005, 2006); Noronha and Shovkovy (2007); Fayazbakhsh and Sadooghi (2010), induce anisotropies in the pairing pattern Yu and Shovkovy (2012), and preferentially stabilize specific superconducting phases Feng et al. (2010). A systematic treatment of neutrino emission in magnetized color-superconducting quark matter is therefore crucial for developing reliable models of the thermal evolution and observable signatures of strongly magnetized compact stars.
@@
CFL color-flavor-locked
2SC two-flavor color-superconducting